From: Flemming Videbaek (videbaek@rcf.rhic.bnl.gov)
Date: Tue Apr 22 2003 - 19:45:41 EDT
Dear Collaborator, The pp startup is making god progress. This week RHIC is still in the startup mode with Beam development for ~ 2 shifts a day and experiment availability for the owl shifts. The Physics running i.e. all time devoted to experiments is expected to start on Monday 28. The Wednesday coordination meeting should clarify this. Last night we took some data with FS at 8 deg low field and MRS at 40 deg. The quality of data looks good, and I believe we are ready for the physics runs, apart from the setup for the polarized luminosity measurements. The plan will be from now until the start of the physics run to have partial manning of the experiment with the aim of completing some of the survey measurements we still need. Tonight has been announced to be for beam-experiments, tomorrow Wednesday is a 12 hour access from ~ 8 am to 8 pm, in which the T3 DC will be re-installed. Tomasz Kozik and Radek have had T3 disassembled, repaired, cleaned and pre-assembled HV and modules ready for mounting an connecting to readout tomorrow. Having T3 back in service will enhance the quality of the FS data. There may be only little beam in the owl shift immeiately following the opening judging the normal pattern after opening, but subsequent nights should be good for data taking at lower fields. I have placed a preliminary shift schedule on the operator web page. Since the expected schedule is really not settled, please make sure to coordinate with me we and Dana what is expected from the machine side on each day. Regards Flemming -- Flemming Videbaek Brookhaven National Laboratory Physics Department tlf 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2.1.5 : Tue Apr 22 2003 - 19:49:26 EDT